OpenStudy (anonymous):

A polynomial P(x) and a divisor d(x) are given. Use long division to find the quotient Q(x) and the remainder R(x) when P(x) is divided by d(x), and express P(x) in the form d(x)*Q(x)+R(x). P(x)=3x^3-x^2+2x+5; d(x)=x+1

myininaya (myininaya):

3x^2-4x+6 x+1 |```````````````````` 3x^3 - x^2+2x+5 -(3x^3+3x^2) --------------- -4x^2+2x+5 -(-4x^2-4x) --------------- 6x+5 -(6x+6) ------ -1

OpenStudy (anonymous):

wow u deserve a trophy

OpenStudy (anonymous):

I know, I was typing it out and then I saw the post and just decided he deserved my vote

myininaya (myininaya):

\[\frac{3x^2-x^2+2x+5}{x+1}=3x^2-4x+6-\frac{1}{x+1}\]

myininaya (myininaya):

im a she lol

OpenStudy (anonymous):

myininaya, you're really good at math

OpenStudy (anonymous):

ok, she. Damn me for my gender stereotyping. She's good, but what's really impressing me is how good she is at working this interface. I could do the math on paper, but it was going to take me about 4 times as long as her to get it all put into the chat....

myininaya (myininaya):

one day i think us females will be able to catch up to you males watch your backs! lol we are evolving (hopefully lol)

OpenStudy (anonymous):

ur living proof right here

myininaya (myininaya):

lol